OpenStudy (anonymous):

How do I prove this trig equation? I have to substitute identities to match the left side of the equation to the right. 1 + sec²x sin²x = sec²x

OpenStudy (anonymous):

If you turn sec^2x into 1/cos^2x, you'd end up with: 1+(sin^2 x)/(cos^2 x) = sec^2 x. Now, sinx/cosx=tanx, so now you have: 1+tan^2 x=sec^2 x, and that is an identity.

OpenStudy (anonymous):

Thanks so much! I was spending way too long on this problem.

OpenStudy (anonymous):

No problem!